OpenStudy (anonymous):

please help Im so lost length of a sides is 3 sqroot 3cm find the area of a square

OpenStudy (anonymous):

It's not as hard as you might think... If you know one side of a square, how do you find area? Easy number example: all sides are 2 cm... So how do you find area for a square with sides 2 cm?

OpenStudy (anonymous):

mulitiply the base times its self

OpenStudy (anonymous):

It's just 2 x 2 = 4 square cm, right? You just multiply Length X Width, but both sides are the same length... it's base x itself, as you said.

OpenStudy (anonymous):

So here, your side length looks "ugly" but it's just a number... 3 sqrt(3) so 3 sqrt(3) X 3 sqrt(3) is the area

OpenStudy (anonymous):

so 27

OpenStudy (anonymous):

yep... \[3\sqrt{3} * 3\sqrt{3} = 3*3*\sqrt{3}*\sqrt{3} = 3*3*(\sqrt{3})^{2} = 3* 3* 3 = 27\]

OpenStudy (anonymous):

It's probably more to teach you how to get comfortable multiplying things with square roots in them than it is to talk about area of a square.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

But you got it :) Good job.